# Quantum memory at nonzero temperature in a thermodynamically trivial system

### v1

@@ Appendix D Open quantum system dynamics
This appendix contains details of our discussion of MFQEC and passive error correction.
D.1 Lindblad dynamics
@@
and as in the main text, the stabilizers $\{S_{a}\}$ are a set of mutually commuting Pauli strings, i.e., $S_{a}^{2}=\mathds{1}$ and $[S_{a},S_{b}]=0$ for all $a$, $b$. In this appendix, we assume that each $S_{a}$ is independent (in the sense that one cannot be built out of a product of any others), as was the case for the non-redundant codes described above. The Gibbs state is diagonal in a basis that consists of eigenvectors of all Pauli strings $\{S_{a}\}$, which we can denote as $\sigma=\sum_{s}\sigma_{s}|s\rangle\langle s|$. Notice that with this choice, there remain $2^{K}$ possible $s$ for each set of stabilizer eigenvalues, and we will be a little lazy and omit that explicit notation.
@@
where $\mathbf{n}\equiv\{n_{a}\}_{a\in\mathcal{A}_{P}}$ and $n_{a}\in\{-1,+1\}$, which defines the projector $\Pi_{P}(\mathbf{n})$. Note that $P=\sum_{\mathbf{n}}P(\mathbf{n})$. Since we have an LDPC CSS code, any local operator $P$ is only involved in a finite number of stabilizers, and each stabilizer has support on a finite number of Pauli operators, so the jump operator $P(\mathbf{n})$ is also $\mathsf{k}$-local. A $\mathsf{k}$-local Lindbladian with stationary state $\sigma$ can then be found explicitly Guo et al. (to appear):
@@
correspoding to “Gibbs sampling” dynamics that protects the Gibbs mixed state.
@@
D.2 Fault tolerance
Now let us relate the Gibbs sampling dynamics above to quantum error correction. For simplicity, we focus on correcting for single-qubit Pauli errors, with generic correction strategies described in Guo et al. (to appear). If $Q$ is a Pauli string acting on O(1) sites, then adding such an error modifies our Lindbladian to:
@@ D.2 Fault tolerance
We will consider an ideal protocol where we apply $Y$ errors deliberately at some rate, to ensure that all $\gamma_{P(\mathbf{n})}>0$, which will shortly prove convenient. Our goal is to show that $\mathcal{L}$ is an exact Gibbs sampler, which will be satisfied so long as the detailed balance condition Guo et al. (to appear)
$(\gamma+\delta\gamma)_{P(\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}=(\gamma+\delta\gamma)_{P(-\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}\mathrm{e}^{-2\beta\sum_{\text{acomm }a}n_{a}}.$ (85)
Here $\mathbf{n}_{\mathrm{acomm/comm}}$ denote the stabilizer values of the anticommuting/commuting stabilizers that act on the same site as single-qubit Pauli $P$. (85) is clearly solved if
$(\gamma+\delta\gamma)_{P(\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}=\gamma_{0}\times\max\left(1,\mathrm{e}^{-2\beta\sum_{\text{acomm }a}n_{a}}\right)=\gamma^{\mathrm{Gibbs}}_{P(\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}$ (86)
@@
D.3 Sort channel implementation

### v2

@@ Appendix E Open quantum system dynamics
This section contains details of our discussion of MFQEC and passive error correction.
E.1 Lindblad dynamics
@@
and as in the main text, the stabilizers $\{S_{a}\}$ are a set of mutually commuting Pauli strings, i.e., $S_{a}^{2}=\mathds{1}$ and $[S_{a},S_{b}]=0$ for all $a$, $b$. In this section, we assume that each $S_{a}$ is independent (in the sense that one cannot be built out of a product of any others), as was the case for the non-redundant codes described above. The Gibbs state is diagonal in a basis that consists of eigenvectors of all Pauli strings $\{S_{a}\}$, which we can denote as $\sigma=\sum_{s}\sigma_{s}|s\rangle\langle s|$. Notice that with this choice, there remain $2^{K}$ possible $s$ for each set of stabilizer eigenvalues, and we will be a little lazy and omit that explicit notation.
@@
where $\mathbf{n}\equiv\{n_{a}\}_{a\in\mathcal{A}_{P}}$ and $n_{a}\in\{-1,+1\}$, which defines the projector $\Pi_{P}(\mathbf{n})$. Note that $P=\sum_{\mathbf{n}}P(\mathbf{n})$. Since we have an LDPC CSS code, any local operator $P$ is only involved in a finite number of stabilizers, and each stabilizer has support on a finite number of Pauli operators, so the jump operator $P(\mathbf{n})$ is also $\mathsf{k}$-local. A $\mathsf{k}$-local Lindbladian with stationary state $\sigma$ can then be found explicitly Guo et al. (2024):
@@
corresponding to “Gibbs sampling” dynamics that protects the Gibbs mixed state.
@@
E.2 Implementing a perfect Gibbs sampler
Let us illustrate how to perform MFQEC for our Gibbs sampler, using $\leq 3$-qubit gates, and reset. For simplicity, we focus only on the step that “corrects” $X$ errors on qubit $i$, as the story is analogous for $Z$ errors and other qubits. For now we also assume that the circuit has no unknown errors; we will relax that assumption shortly. Let ancilla qubits $a_{1},\ldots,a_{\Delta_{C}}$ store the “measurement outcomes” for the $\Delta_{C}$ $Z$-checks $C_{1},\dots,C_{q}$ containing physical qubit $i$. The first step of MFQEC is to apply the unitary
where $\mathsf{CNOT}_{i\rightarrow j}=\frac{1+Z_{i}}{2}+\frac{1-Z_{i}}{2}X_{j}$. In the second step, we apply the $\mathsf{SORT}_{\Delta_{C}}$ channel, which shuffles the ancilla qubits such that the zeros are to the left and the ones are to the right. We later provide an implementation for $\mathsf{SORT}$. In order to realize a perfect Gibbs sampler, we apply a coherent gate that implements the following controlled-unitary using one additional ancilla qubit $c$, which is again initialized in state $|0\rangle$:
where we choose each angle $\theta_{j}$ such that
E.3 Fault tolerance
Now let us relate the Gibbs sampling dynamics above to quantum error correction. For simplicity, we focus on correcting single-qubit Pauli errors, with generic correction strategies described in Guo et al. (2024). If $Q$ is a Pauli string acting on O(1) sites, then adding such an error modifies our Lindbladian to:
@@ E.3 Fault tolerance
We will consider an ideal protocol where we apply $Y$ errors deliberately at some rate, to ensure that all $\gamma_{P(\mathbf{n})}>0$, which will shortly prove convenient. Our goal is to show that $\mathcal{L}$ is an exact Gibbs sampler, which will be satisfied so long as the detailed balance condition Guo et al. (2024)
$(\gamma+\delta\gamma)_{P(\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}=(\gamma+\delta\gamma)_{P(-\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}\mathrm{e}^{-2\beta\sum_{\text{acomm }a}n_{a}}$ (98)
holds. Here $\mathbf{n}_{\mathrm{acomm/comm}}$ denote the stabilizer values of the anticommuting/commuting stabilizers that act on the same site as single-qubit Pauli $P$. (98) is clearly solved if
$(\gamma+\delta\gamma)_{P(\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}=\gamma_{0}\times\max\left(1,\mathrm{e}^{-2\beta\sum_{\text{acomm }a}n_{a}}\right)=\gamma^{\mathrm{Gibbs}}_{P(\mathbf{n}_{\mathrm{acomm}},\mathbf{n}_{\mathrm{comm}})}\,,$ (99)
@@
E.4 Sort channel implementation